# CASH: Capability-Aware Shared Hypernetworks for Flexible Heterogeneous Multi-Robot Coordination

### v4

@@ Abstract
**footnotetext: Equal Contribution
@@ 2 Related Works
Architectures for Heterogeneous Coordination: Shared-parameter architectures – where all agents share a single set of learnable parameters – improve cooperation, learning efficiency, and scalability [5, 6]. However, sharing parameters can prohibit diverse agent behaviors and unique roles. While this limitation is immaterial for homogeneous teams (e.g., formation control [7] and path planning [8, 9]), it can critically limit heterogeneous teams. To extend shared-parameter architectures to heterogeneous teams, prior work has simply appended unique information about each agent to the input [10, 11]; typically, a unique ID assigned to each agent. While such ID-based designs tend to improve efficiency and scalability, recent work has demonstrated that they often fail to learn sufficient behavioral diversity and are not robust to noisy environments [12, 1].
@@ 3 Problem Formulation and Objectives
We specifically focus on teams where the robots have heterogeneous capabilities (e.g., speed and sensing radius) [19, 11]. To enable a richer specification of robot heterogeneity, we use $C^{t}=\{c^{t}_{1},...,c^{t}_{n}\}$ to denote the team’s capabilities, where $c^{t}_{i}\in\mathbb{R}^{m}$ represents the capabilities of $i$th robot at time $t$. Therefore, we modify the Dec-POMDP definition by giving each policy $\pi_{i}(\cdot)$ access to the team’s collective capabilities, yielding $a^{t}_{i}\sim\pi_{i}(o^{t}_{i},C^{t})$.
@@ 5.1 Experiments on JaxMARL
    RNN-IMP: The standard RNN-based architecture that is implicitly conditioned on capabilities by way of processing observations. Indeed, some capabilities (e.g. speed) can be inferred solely from observations over time. This baseline questions the need for explicit conditioning on capabilities.
@@ 5.2 Experiments on the Robotarium
CASH adapts to online changes to robot capabilities. We investigated how CASH handles online changes to robot capabilities by taking the trained policies for each task and zero-shot deploying them in the Failure and Battery Drain scenarios. As expected, we observe a drop in task performance across all baselines in Tab. 3. However, CASH consistently outperforms the baselines when dealing with the challenging unplanned changes to robot dynamics and observations during policy execution. This is likely due to CASH’s better capability grounding that allows it to dynamically adjust the robots’ behaviors even during a rollout, a critical feature to enable robustness in learned heterogeneous teaming.
@@ 7 Limitations
Another limitation is that the JaxMARL tasks which we evaluated CASH on are simpler than the most complex tasks in heterogeneous multi-agent coordination. For instance, our Firefighting task can be seen as a simplified version of a firefighting simulator like FireCommander [13]. However, our results show that despite the simplicity of these tasks, they have the crucial property that better grounding of heterogeneous capabilities tends to yield better performance, as evidenced by RNN-EXP outperforming RNN-IMP and CASH outperforming RNN-EXP. In our opinion, this is the most important characteristic relevant to heterogeneous robots, and abstracting away other details (e.g. more realistic communication, continuous action spaces, more dynamic environments) allowed us to tease out how policies can more effectively capture the relationship between robot capabilities and behaviors. Further, our experiments in the Robotarium show that the benefits of CASH observed in simulation hold for real multi-robot teams as well.
@@ Acknowledgments
If a paper is accepted, the final camera-ready version will (and probably should) include acknowledgments. All acknowledgments go at the end of the paper, including thanks to reviewers who gave useful comments, to colleagues who contributed to the ideas, and to funding agencies and corporate sponsors that provided financial support.
@@ Appendix A Evaluating Behavioral Diversity with SND
Figure 5: Impact of removing Layer Normalization from the Hyper Adapter of CASH when training with three learning paradigms across two simulation tasks. It is evident that LayerNorm is a crucial component in stabilizing the training of the hypernetwork within CASH.
@@ Appendix B Layer normalization is crucial to CASH
To show the importance of adding layer normalization (LayerNorm) [36] when training hypernetworks for multi-agent coordination problems, we present an ablation over the inclusion of LayerNorm in CASH’s Hyper Adapter in Figure 5. Across all tasks and learning algorithms, the inclusion of LayerNorm improves training stability and decreases variance between seeds. In all cases except one (MAPPO / Mining), it results in a significant improvement on the final converged returns. DAgger is a notable outlier in terms of performance gap with and without LayerNorm; we see that when training with DAgger, the performance of CASH without LayerNorm in Mining plummets around timestep 4000 and never recovers, while in Firefighting, CASH without LayerNorm never demonstrates better than random performance.
Taken together, these results show that layer normalization improves training stability for hypernetworks, and that this improved stability is crucial for our architecture. We note that adding layer normalization to hypernetworks was suggested by the original Hypernetworks paper [2] to improve gradient flow, which our findings here corroborate. Surprisingly, prior work on using hypernetworks in meta-RL for generalization to new tasks [27, 37] does not include LayerNorm in the hypernetwork, even though we found it to be critically important for CASH. We speculate that these works are able to achieve success despite omitting normalization from their hypernetworks because the input to their hypernetworks is a strong pretrained encoder. By contrast, our hypernetwork is conditioned directly on task observations and team capabilities, without the benefit of an encoder for preprocessing. As a result, we found that a deeper four-layer hypernetwork was necessary. We leave a more thorough investigation of normalization schemes for training deep hypernetworks to future work.
@@ C.1 Hyper-Adapter
Our experiments revealed that a deeper four-layer hypernetwork is better for our use case. However, consistent with prior work using hypernetworks in RL [27], we found it difficult to optimize a standard four-layer hypernetwork, both in reinforcement and imitation learning regimes. We believe this is due to the combination of instabilities inherent in both the learning paradigms as well as hypernetworks themselves. However, we found that simply adding layer normalization (LayerNorm) [36] before each ReLU activation in our hypernetwork stabilizes learning and greatly improves performance (see our supplementary material for ablation experiments and further discussion). This finding aligns with recent work suggesting LayerNorm can improve training stability for deeper networks across a wide range of RL domains [38, 39].
For CASH, we implemented the Hyper Adapter as two separate hypernetworks, one which generates the weights of the target linear layer and the other which generates the biases. We found that initializing both hypernetwork with orthogonal weights, zero-bias was conducive to good task success, which is common in other works. However, we noticed a slight performance increase when setting the scale of those weights to be 0 for the bias-generating hypernet and 0.2 for weight-generating hypernet. We could not find a reference for this odd initialization scheme in the literature. For INDV, we condition each robot’s policy on observations, the ego robot’s capabilities, and the capabilities of its teammates. This is analogous the conditioning of RNN-EXP.
@@ D.1 Firefighting
In Tab. 9 and Tab. 8 we detail our randomly selected train and test teams for Firefighting and Mining. More details on how we chose these teams and how they were used can be found in the main body of Chapter 3. For context, in Firefighting the fires range from 0.2-0.3 strength and the training distribution of agents ranged between 0.1-0.3 firefighting capacity and 1-3 acceleration. For testing teams, we sample agents from out of bounds in both firefighting and acceleration, but match other agents to ensure task feasibility.
@@
Table 8: Firefighting Training Teams
@@
Table 9: Firefighting Testing Teams
@@ D.2 Mining
Table 10: Mining Training Teams
@@
Table 11: Mining Testing Teams
@@ D.3 Material Transport
Table 12: Material Transport Capability Sets
@@
Table 13: Material Transport Deployment Teams
@@ D.4 Predator Capture Prey
Table 14: Predator Capture Prey Capability Sets
@@
Table 15: Predator Capture Prey Deployment Teams
@@ Appendix G Additional JaxMARL Results
G.1 Training on Larger Teams
@@ G.1.3 Results
Table 19: 12 agent Firefighting Training Teams
@@
Table 20: 12 agent Firefighting Testing Teams
@@
Table 21: 12 agent Mining Training Teams
@@
Table 22: 12 agent Mining Testing Teams

### v5

@@ Abstract
**footnotetext: Equal Contributionfootnotetext: Project Website: https://star-lab.cc.gatech.edu/papers/fu-jain-CASH-CoRL/
@@ 2 Related Works
Architectures for Heterogeneous Coordination: Shared-parameter architectures – where all agents share a single set of learnable parameters – improve cooperation, learning efficiency, and scalability [5, 6]. However, sharing parameters can prohibit diverse agent behaviors and unique roles. While this limitation is often immaterial for homogeneous teams (e.g., formation control [7] and path planning [8, 9]), it can critically limit heterogeneous teams. To extend shared-parameter architectures to heterogeneous teams, prior work has simply appended unique information about each agent to the input [10, 11]; typically, a unique ID assigned to each agent. While such ID-based designs tend to improve efficiency and scalability, recent work has demonstrated that they often fail to learn sufficient behavioral diversity and are not robust to noisy environments [12, 1].
@@ 3 Problem Formulation and Objectives
We specifically focus on teams where the robots have heterogeneous capabilities (e.g., speed and sensing radius) [19, 11]; such nominal capabilities as we define them can be readily obtained from robot specifications or sensors. To enable a richer specification of robot heterogeneity, we use $C^{t}=\{c^{t}_{1},...,c^{t}_{n}\}$ to denote the team’s capabilities, where $c^{t}_{i}\in\mathbb{R}^{m}$ represents the capabilities of $i$th robot at time $t$. Therefore, we modify the Dec-POMDP definition by giving each policy $\pi_{i}(\cdot)$ access to the team’s collective capabilities, yielding $a^{t}_{i}\sim\pi_{i}(o^{t}_{i},C^{t})$.
@@ 5.1 Experiments on JaxMARL
    RNN-IMP: The standard RNN-based architecture that is implicitly conditioned on capabilities by way of processing observations. Indeed, some capabilities (e.g. speed) can be inferred solely from observations over time with memory-enabled architectures (e.g., GRU). This baseline questions the need for explicit conditioning on capabilities.
@@ 5.2 Experiments on the Robotarium
CASH adapts to online changes to robot capabilities. We investigated how CASH handles online changes to robot capabilities by taking the trained policies for each task and zero-shot deploying them in the Failure and Battery Drain scenarios. As expected, we observe a drop in task performance across all baselines in Tab. 3. However, CASH consistently outperforms the baselines when dealing with the challenging unplanned changes to robot capabilities and observations during policy execution. This is likely due to CASH’s better capability grounding that allows it to dynamically adjust the robots’ behaviors even during a rollout, a critical feature to enable robustness in learned heterogeneous teaming.
@@ 7 Limitations
A constraint of CASH is that it cannot handle entirely new capabilities dimensions during inference. This is a natural limitation of most neural network learning paradigms with the exception of large reasoning models that leverage massive amounts of data and tokenization, which could potentially handle new capabilities during inference time. Furthermore, it is fair to assume for most single-domain policies that the capability dimensions are the same for training and inference.
Another limitation is that the JaxMARL tasks which we evaluated CASH on are simpler than the most complex tasks in heterogeneous multi-agent coordination. For instance, our Firefighting task can be seen as a simplified version of a firefighting simulator like FireCommander [13]. However, our results show that despite the simplicity of these tasks, they have the crucial property that better grounding of heterogeneous capabilities tends to yield better performance, as evidenced by RNN-EXP outperforming RNN-IMP and CASH outperforming RNN-EXP. In our opinion, this is the most important characteristic relevant to heterogeneous robots, and abstracting away other details (e.g. more realistic communication, continuous action spaces, sensing and actuation noise, more dynamic environments) allowed us to tease out how policies can more effectively capture the relationship between robot capabilities and behaviors. Further, our experiments in the Robotarium show that the benefits of CASH observed in simulation hold for real multi-robot teams as well.
@@ Acknowledgments
The authors would like to thank the reviewers for their detailed and constructive feedback to help strengthen the paper. We would also like to thank the Robotarium by Georgia Tech for it’s accessibility to multi-robot hardware for real-robot experimentation and testing. This work was supported in part by the Army Research Lab under Grant W911NF17-2-0181 (DCIST CRA).
@@ Appendix A Evaluating Behavioral Diversity with SND
Figure 5: This figure shows the impact on training returns as a result of removing Layer Normalization from the Hyper Adapter of CASH. The results are presented for three learning paradigms across two simulation tasks. It is evident that LayerNorm is a crucial component in stabilizing the training of the hypernetwork within CASH.
@@ Appendix B Layer normalization is crucial to CASH
To show the importance of adding layer normalization (LayerNorm) [34] when training hypernetworks for multi-agent coordination problems, we present an ablation over the inclusion of LayerNorm in CASH’s Hyper Adapter in Figure 5. Across all tasks and learning algorithms, the inclusion of LayerNorm improves training stability and decreases variance between seeds, which aligns with recent works suggesting the benefit of LayerNorm for improving training stability of deeper networks in RL [35, 36]. In all cases except one (MAPPO / Mining), it results in a significant improvement on the final converged returns. DAgger is a notable outlier in terms of performance gap with and without LayerNorm; we see that when training with DAgger, the performance of CASH without LayerNorm in Mining plummets around timestep 4000 and never recovers, while in Firefighting, CASH without LayerNorm never demonstrates better than random performance.
Taken together, these results show that layer normalization improves training stability for hypernetworks, and that this improved stability is crucial for our architecture. We note that adding layer normalization to hypernetworks was suggested by the original Hypernetworks paper [2] to improve gradient flow, which our findings corroborate. Surprisingly, prior work on using hypernetworks in meta-RL for generalization to new tasks [27, 37] does not include LayerNorm in the hypernetwork, even though we found it to be critically important for CASH. We speculate that these works are able to achieve success despite omitting normalization from their hypernetworks because the input to their hypernetworks is a strong pretrained encoder. By contrast, our hypernetwork is conditioned directly on task observations and team capabilities, without the benefit of an encoder for preprocessing. We leave a more thorough investigation of normalization schemes for training deep hypernetworks to future work.
@@ C.1 Hyper-Adapter
Our experiments revealed that a deeper four-layer hypernetwork is better for our use case. However, consistent with prior work using hypernetworks in RL [27], we found it difficult to optimize a standard four-layer hypernetwork, both in reinforcement and imitation learning regimes. We believe this is due to the combination of instabilities inherent in both the learning paradigms as well as hypernetworks themselves. However, we found that simply adding layer normalization (LayerNorm) [34] before each ReLU activation in our hypernetwork stabilizes learning and greatly improves performance (see Appendix B for more details).
For CASH, we implemented the Hyper Adapter as two separate hypernetworks, one which generates the weights of the target linear layer and the other which generates the biases. We found that initializing both hypernetwork with orthogonal weights, zero-bias was conducive to good task success, which is common in other works. However, we noticed a slight performance increase when setting the scale of those weights to be 0 for the bias-generating hypernet and 0.2 for weight-generating hypernet. We could not find a reference for this odd initialization scheme in the literature.
@@ D.1 Firefighting
In Table 8 and Table 9 we detail our randomly selected train and test teams for Firefighting. More details on how we chose these teams and how they were used can be found in the main body of Chapter 3. For context, in Firefighting the fires range from 0.2-0.3 strength and the training distribution of agents ranged between 0.1-0.3 firefighting capacity and 1-3 acceleration. For testing teams, we sample agents from out of bounds in both radius and acceleration, but match other agents to ensure task feasibility.
@@
Table 8: Firefighting training robots.
@@
Table 9: Firefighting testing teams.
@@ D.2 Mining
Table 10: Mining training teams.
@@
Table 11: Mining testing teams.
@@ D.3 Material Transport
Table 12: Material transport capability sets.
@@
Table 13: Material transport deployment teams.
@@ D.4 Predator Capture Prey
Table 14: Predator capture prey capability sets.
@@
Table 15: Predator capture prey deployment teams.
@@ Appendix G Additional JaxMARL Results
G.1 Comparison against ID-based methods
To further contextualize the INDV results, we compare against shared-parameter ID-based methods where a unique one-hot ID is appended to the observation space of each agent. ID-based methods represent a shared parameter baseline that cannot generalize to unseen robots. The experiment setup is identical to that in Section 2, but includes results from an agent ID baseline labeled RNN-ID. We plot the returns and SND for the Firefighting and Mining tasks in Figure 9. Expectedly, the shared encoder does improve sample efficiency when compared to individualized policies, and RNN-ID learns a lower level of behavioral diversity.
Figure 9: ID-based methods are more sample efficient than individualized policies (see returns), but learn a lower level of behavioral diversity (see SND). CASH matches or outperforms ID-based methods in both sample efficiency and behavioral diversity.
G.2 Training on Larger Teams
@@ G.2.3 Results
Table 19: 12 agent Firefighting training teams.
@@
Table 20: 12 agent Firefighting testing teams.
@@
Table 21: 12 agent Mining training teams.
@@
Table 22: 12 agent Mining testing teams.